Swathmore Clothing Corporation grants its customers 30 days’ credit. The company uses the allowance method for its uncollectible accounts receivable. During the year, a monthly bad debt accrual is made by multiplying 2% times the amount of credit sales for the month. At the fiscal year-end of December 31, an aging of accounts receivable schedule is prepared and the allowance for uncollectible accounts is adjusted accordingly.

At the end of 2020, accounts receivable were $608,000 and the allowance account had a credit balance of $72,000. Accounts receivable activity for 2021 was as follows:

Beginning balance Credit sales $ 608,000 2,790,000 (2,653,000) (56,000) Collections Write-offs 689,000 Ending balance The comRequired:
1. Prepare a summary journal entry to record the monthly bad debt accrual and the write-offs during the year.
2. Prepare the necessary year-end adjusting entry for bad debt expense.
3-a. What is total bad debt expense for 2021?
3-b. How would accounts receivable appear in the 2021 balance sheet?
Answer #1

SOLUTION

1.

S.No. Accounts titles and Explanation Debit ($) Credit ($)
1. Bad debt expense 55,800
Allowance for uncollectible accounts 55,800
2. Allowance for uncollectible accounts 56,000
Accounts receivable 56,000

Bad debt expense = Credit sales * 2%

= 2,790,000 * 2% = 55,800

2. Journal entry-

Accounts titles and Explanation Debit ($) Credit ($)
Bad debt expense 14,890
  Allowance for uncollectible accounts 14,890

Working-

Estimated uncollectible accounts-

Details Amount ($) (A) Percentage uncollectible (B) Estimated uncollectible amounts (A*B)
0-60 days 455,000 5% 22,750
61-90 days 76,000 14% 10,640
91-120 days 66,000 25% 16,500
Over 120 days 92,000 40% 36,800
689,000 86,690

Bad debt expense at the year end-

Allowance for uncollectible accounts:
Beginning balance 72,000
Add: Monthly bad debt accruals 55,800
Deduct: Write offs (56,000)
Balance before year end adjustment 71,800
Required allowance 86,690
Required year end increase in allowance 14,890

3A. Total bad debts = Bad debts during the year + Year end Bad debt expense

= 55,800 + 14,890 = 70,690

3B.

Assets Amount ($)
Current assets:
Accounts receivable (net) 602,310

Accounts receivable (net) = Accounts receivable (gross) - Allowance for uncollectible account

= 689,000 - 86,690 = 602,310